Circular No. 5643

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

EU 1737–132 S. Brandt, A. J. Castro-Tirado, and N. Lund report: “A new x-ray

transient in Serpens Cauda has been discovered by the WATCH wide-field monitor on the ESA EURECA satellite. The transient was first observed on Oct. 11, at α = 17^{h}37^{m}20^{s}, δ = −13^{o}15′.0 (equinox 1950.0; error radius 45′). The source brightness at the time of discovery was ∼ 500 mCrab (above 6 keV), and remained at about this level at least until Oct. 18. In the spectral band between 6 and 15 keV, the source is now only slightly harder than that of Sco X-1, though during the first two days of the outburst the spectrum was harder. Our limited sensitivity does not allow us to detect the source above 15 keV. The source may be similar to that of dwarf x-ray transients such as N Mus 1991 and GRO J0422+32. Follow up at other wavelengths is encouraged.”

PERIODIC COMET SWIFT-TUTTLE (1992t) P. Colom, D. Bockelée-Morvan, G. Bourgois, J. Crovisier, E. Gérard,

and L. Jorda, Paris-Meudon Observatory, communicate: “The main lines of the OH radical at 1667 and 1665 MHz were detected in emission in P/Swift-Tuttle on Oct. 15 with the Nançay Radio Telescope. For the period Oct. 16-20, the 1667-MHz profile has a peak flux density of 86 ± 14 mJy, a FWHM of 2 km/s, and is blueshifted by 0.3 km/s with respect to the radial velocity of the nucleus. The line area of 184 ± 17 mJy km/s corresponds to an OH production rate of 8 × 10^{28} molecules/s.”

J. V. Scotti, University of Arizona, reports that CCD images were obtained of this comet on Oct. 18.1, 19.1, and 21.1 UT with the Spacewatch telescope, showing an ion tail extending at least 15′ toward the northnortheast. There was a noticeable increase in inner coma activity, as well as in overall brightness, between Oct. 18.1 and 19.1. On Oct. 21.097 UT, there was a 10′.5 coma surrounding an inner coma that is asymmetrical toward the west (though the overall coma is nearly circular). A fan-shaped jet to the west dissipated slightly between Oct. 18 and 21, and the tail was a bit fainter on Oct. 21 than on previous nights. The fan-shaped jet structure on Oct. 21 runs from p.a. 242^{o} to 313^{o}; the main tail extends toward p.a. 24^{o}, and there are two faint streamers bracketing the tail at p.a. 9^{o} and 50^{o}, each extending ∼ 5′ and perhaps further.

1992 October 22 Daniel W. E. Green